OpenStudy (anonymous):

6. Use factoring to solve the quadratic equations below, if possible. a. 2x2 + 3x - 54 = 0 b. 4w2 + 6 =11w c. 3t2 -13t =10 d. 2x(4x + 3) = 3+ x e. 18z2 + 21z = 4(z -1) f. 8 -13p = 6 p(6 - p) g. 24q2 = 4q + 8

OpenStudy (anonymous):

Which parts specifically are you having trouble with?

OpenStudy (anonymous):

mmm i will do no:b then u will do the rest 4w^2+6=11w then take 11w to the left side. 4w^2+6-11w=0 (it changes sign since it is passing over an equal sing) 4w^2-11w+6=0 then you find the factors (4w-3)(w-2)=0 then w=3/4 ; w=2